OpenStudy (anonymous):

What is the magnitude of the electric field that will balance the weight of an electron? The acceleration of gravity is 9.81 m/s2 .

OpenStudy (anonymous):

E=F/q e charge = 1.6*10^-(19) F=mg = (9.1*10^(-31))kg*9.81m/s^2

OpenStudy (anonymous):

thx for the help, but it said i was wrong =(